40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 107     Entry time: Wed Apr 10 00:40:30 2013
Author: Zach 
Type: Optics 
Category: Configuration 
Subject: fauxMC locked 

[Koji, Zach]

Tonight, we locked the "fauxMC". We obtained a visibility of >99%.

Koji had aligned it roughly last night, but we wanted to have a couple steering mirrors in the path for this practice cavity (the periscope mirrors will serve this function in the real setup), so we marked the alignment with irises and installed two extra mirrors.

After obtaining flashes with the WinCam placed at the output coupler, we removed the WinCam and put a CCD camera at one of the curved mirror transmissions and used this to get a strong TEM00 flash. Then, we installed the REFL PD/CCD, swept the laser PZT and optimized the alignment by minimizing the REFL dips. Finally, we connected the RF electronics and locked the cavity with the LB box. We used whatever cables we had around to trim the RF phase, and then Koji made some nice SMA cables at the 40m.

One thing we noticed was that we don't have enough actuation range to keep the cavity locked for very long---even with the HV amp (100V). We are going to offload to the NPRO temperature using an SR560 or pomona box circuit. We may also make an enclosure for the cavity to protect it from the HEPA blasting.

Tomorrow, after we do the above things, we will practice measuring the transmission, length (FSR) and mode spectrum of the cavity before moving on to the real McCoy.

fauxMC_labeled.jpg

ELOG V3.1.3-